An extra-large pizza is circular with a diameter of 12 inches. Pizza sauce is spread on the pizza dough at a rate of 1.1 mL of pizza sauce per square inch (in2) of crust. How much pizza sauce (rounded to the nearest 1 mL) is required to cover the entire pizza, if a 1-inch crust is to be left around the edge of the entire pizza?


1
Expert's answer
2022-04-21T04:35:04-0400

Square pizza:


"S_{p}=\frac{\pi d^{2}}{4}\\"


"d=12 - 1- 1 = 10(in)" - diameter of pizza slice covered with sauce.


"S_{p_s}=\frac{\pi 10^{2}}{4}=78.539 (in^{2})"


"78.539 * 1.1 = 86.3929 \approx87 (ml)" - sauce need.